OpenStudy (anonymous):

A proportional relationship between the number of pounds of cabbage (x) and the price in dollars (y) is graphed, and the ordered pair (5, 2) is on the graphed line. Part A: what is the price of 1 pound of cabbage? Show your work, including The proportion you use to determine the price. Part B: What does the ordered pair (10, 4) on the graph represent? Explain in words.

OpenStudy (danjs):

Proportional relationship y = k*x , where k is a constant

OpenStudy (danjs):

(x,y)=(5,2) k = y/x = 2/5 so, y=(2/5)*x

OpenStudy (danjs):

what is y when x is 1 pound of cabbage? y=(2/5)*1 = 2/5

OpenStudy (danjs):

(10,4) , 4 lbs of cabbage costs 10 dollars